OpenStudy (anonymous):

simplify and state excluded values. 2a+2/a^2-1

OpenStudy (mertsj):

\[\frac{2a+2}{a^2-1}=\frac{2(a+1)}{(a-1)(a+1)}=\frac{2}{a-1}\]

OpenStudy (mertsj):

a cannot be 1 or -1